OpenStudy (anonymous):

A bike to do tricks with has 20-inch-diameter wheels. Find the area of each wheel. Round. Choose - 1. 314.0 in.2 2. 314.2 in.2 3. 1256.4 in.2 4. 1256.6 in.2

OpenStudy (anonymous):

pi times the radius squared.

OpenStudy (saifoo.khan):

\[Area = \pi r^2\]\[Area = \pi 10^2\]\[Area = option 2.\]

OpenStudy (anonymous):

thank you <3

OpenStudy (saifoo.khan):

No problem! ^_^